wasp 43 ch2 2s subarray
- All 64 frames in one bcd, ~2.7 hours into the observation, have elevated flux across the whole frame (multiplied by ~1E5). It looks like a normal image when displaying with ds9, only the photometry indicates something is wrong. The flux levels in the bcd stack before and after are fine.
- The imask implies the whole image is non-linear
- The uncertainties (_bunc.fits) image is all nan's.
The source of this problem is a corrupted header. The barrel shift and some other header keywords were corrupted in transmission, causing the pipeline to not be able to correctly process the image.